What would be the vertex of the following:
f(x)=x^2+10x+9

Answer by Alan3354(69443)

What would be the vertex of the following:
f(x)=x^2+10x+9
-----------------
The line of symmetry is x = -b/2a
x = -10/2 = -5
The vertex is (-5,f(-5))
= (-5,-16)